OpenStudy (anonymous):

a tree cats a shadow that is 14 feet long at the same time that a near by 2 foot tall pole caSTS A SHADOW that is 1.75 feet long. how tall is the tree?

OpenStudy (anonymous):

Let x be the tree's actual height x:14=2:1.75 14*2=1.75x 16=x

OpenStudy (anonymous):

thank you!(:

OpenStudy (anonymous):

yoour welcome :)

OpenStudy (anonymous):

help me with one more?

OpenStudy (anonymous):

find the simple interest paid for 10 years on a $10,000 loan at 9% per year?

OpenStudy (anonymous):

I=Prt Principal=10000 rate=9% or 0.09 time=10 years I=Prt I=10000*0.09*10 I=9000

OpenStudy (anonymous):

life saver!(:

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

anytime :)